A triangle is placed in a semicircle with a radius of 2mm what is the area of the shaded region use the value 3.14 and do not round the number





1
Expert's answer
2021-02-02T04:37:53-0500

1) with green checkmarks the segments, the area of which (S) I am looking for

2) S of semicircle = "S_s"




S = "S_s" - "S\vartriangle(ABC)"


"S_s = \tfrac{1}{2}\pi*R^2 => \tfrac{1}{2}*3.14*4 mm^2 = 6.28 mm^2"

"S\vartriangle(ABC) = \tfrac{1}{2} BO*AC = \tfrac{1}{2} *2mm*4mm = 4mm^2"

"S = 6.28 - 4 = 2.28 mm^2"